OpenStudy (anonymous):

Solve and check for extraneous solutions 16=3(x-1)-(x-7)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

not sure what "extraneous" solutions means in this context it is a linear equation right? no radials, no denominators

OpenStudy (anonymous):

distribute first on the right both the 3 and the minus sign

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

Ok, so I got 2x+6=16

OpenStudy (anonymous):

think there is a mistake there the \(2x\) is right, but where did the \(6\) come from?

OpenStudy (anonymous):

-1+7

OpenStudy (anonymous):

\[3(x-1)-(x-7)=3x-3-x+7\] is a start

OpenStudy (anonymous):

oops forgot the distributive law didn't you?

OpenStudy (anonymous):

\[3(x-1)=3x-3\times 1\]

OpenStudy (anonymous):

Oh okay! So 2x+4?

OpenStudy (anonymous):

right \[2x+4=16\] takes only 2 steps to solve

OpenStudy (anonymous):

a) subtract 4 b) divide by 2

OpenStudy (anonymous):

2x=10 x=2

OpenStudy (anonymous):

???

OpenStudy (anonymous):

first off \(16-4\neq 10\) and secondly \(\frac{10}{2}\neq 2\) so both steps are wrong

OpenStudy (anonymous):

what is \(16-4\)?

OpenStudy (anonymous):

16-4=12

OpenStudy (anonymous):

ok so \[2x+4=16\\ 2x=12\] is a the first step

OpenStudy (anonymous):

x=6! So sorry, it's late where I am lol so I have a sleepy brain I suppose

OpenStudy (anonymous):

second step is \(x=\frac{12}{2}\)

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

right

OpenStudy (anonymous):

How do we check for extraneous solutions?

OpenStudy (anonymous):

there are no extraneous solutions with linear equations, nothing to check

OpenStudy (anonymous):

you can replace x by 6 in the original equation and see that it works if you like but it is right, nothing is extraneous here

OpenStudy (anonymous):

ok:) Can I ask another? I'll be better haha

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

x/3=x/2-2 Check for extraneous solutions

OpenStudy (anonymous):

is this \[\frac{x}{3}=\frac{x}{2}-2\] like that ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

you have a choice you can work with the fractions or you can get rid of them

OpenStudy (anonymous):

you pick

OpenStudy (anonymous):

get rid of them.

OpenStudy (anonymous):

Actually, which ever is simpler

OpenStudy (anonymous):

then since the least common multiple of 2 and 3 is 6, if you multiply both sides by 6 they will be gone most people i think like getting rid of fractions

OpenStudy (anonymous):

\[6\times \frac{x}{3}=6\times (\frac{x}{2}-2)\]

OpenStudy (anonymous):

I think I did it wrong bc I got 2x/6-3x/6

OpenStudy (anonymous):

since \(\frac{6}{2}=3\) and \(\frac{x}{3}=2\) you end up with \[2x=3x-18\]

OpenStudy (anonymous):

you forgot to get rid of the 6 when you cancelled

OpenStudy (anonymous):

oh okay:)

OpenStudy (anonymous):

\[6\times \frac{x}{3}=2x\] not \(\frac{2x}{6}\)

OpenStudy (anonymous):

so now you have \[2x=3x-18\] which takes 2 steps to solve you see where the \(-18\) came from?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

x=18

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So it wouldnt be extraneous?

OpenStudy (anonymous):

no it would not